A 2.0 mm thick sheet of shielding allows 80% of incident gamma radiation to pass through it.
How many layers of this shielding are required such that less than 30% of incident gamma rays to pass through all layers? (to 2 s.f)
1
Expert's answer
2020-05-28T13:38:23-0400

We know that one layer allows 80% of incident radiation to pass and absorbs 100%-80% = 20% of radiation.

Let "X" be the initial amount of radiation. After passing through one layer, 80% of "X" (or "0.8X" ) will remain. This amount of radiation falls onto the second layer, and after passing through it "0.8\cdot0.8 X = 0.64X" will remain. Therefore, the amount of radiation remaining after passing through "N" layers is "0.8^NX." Let us determine "N" for which "0.8^NX \le 0.3X" or "0.8^N\le0.3" .

"0.8^5 = 0.32768 > 0.3, \;\; 0.8^6 = 0.262144 < 0.3." Therefore, the minimal number of layers is 6 and the total thickness will be "2.0\cdot6 = 1.2\,\mathrm{mm}."